# Insertion space in repulsive active matter

### v1

@@ Abstract
where $\beta=(k_{B}T)^{-1}$ is the inverse temperature, $P$ is the bulk pressure, $V$ is the total volume of the space, $N$ is the number of constituents, $\sigma$ is the particle diameter (or repulsion range), $d$ is the spatial dimension, $\Lambda$ is the thermal de Broglie wavelength, $\langle V_{I}\rangle$ is the averaged (over all values in all configurations) total insertion volume, and $\langle S_{I}\rangle$ its averaged concomitant surface. The insertion volume is expressed in terms of the averaged number of cavities and cavity size as $\langle V_{I}\rangle=\langle N_{\Delta}\rangle\langle\Delta\rangle$ [15, 17].
@@
Figure 2: Insertion space (IS) for repulsive active particles in $d=1$. a) Averaged normalized cavity size as a function of packing fraction $\eta$, as determined from off-lattice (filled points) and on-lattice (unfilled points) simulations of RTPs at various $K$. Analytical results are solid lines; for $K\leq 10.0$ the lines are the weak activity mean-field expressions while for $K>10.0$ we show the high-activity expressions (9-11). Exact passive expressions are from (7). b) The averaged total insertion volume divided by the system volume. c) Averaged number of insertion cavities per particle as a function of $\eta$. d) The geometric compressibility, $Z_{G}$, as a function of $\eta$. Inset is the geometric chemical potential $\mu_{G}$ against $\eta$. e) - h) The same quantities as for a)-d) but for the AOUP (off-lattice) model. High-activity analytics (for $\tau=1000$) are (9-11) with $K=\sqrt{D^{\prime}\tau}/\sigma$. Parameters (RTPs): $N=1024$, $\mu=1$ (off-lattice), $\epsilon=10$ (off-lattice), and $\sigma=1$; AOUPs: $N=1024$, $\mu=1$, $\epsilon=10$, $D^{\prime}=\frac{1}{2}$ and $\sigma=1$.
Firstly, in $d=1$, the non-dimensionlized averaged cavity size $\left\langle\Delta\right\rangle/\sigma\equiv\langle\tilde{\Delta}\rangle$ is the free (unoccupied) space divided by the number of non-touching particle pairs minus one and reads as:
@@
where $P_{C}$ is the probability of two particles being in contact and $\eta=N\sigma/V$ is the packing fraction. In one dimension, $\langle S_{I}\rangle=2$ exactly, implying that the geometric compressibility $Z_{G}$ (1) can be simplified as [13]
@@
with $\langle V_{I}\rangle=V-N2\sigma$ in the dilute case, where $2\sigma$ is the volume excluded by one particle, and $\langle V_{I}\rangle=0$ at full packing. Note, that $P_{C}$ is absent in the original expressions of these quantities [13]. The equilibrium one-dimensional hard-sphere model is the Tonks gas, which is exactly solvable [8], where the IS quantities are given exactly as [14, 15]:
@@
For the active model we consider repulsive run-and-tumble particles (RTPs) in $d=1$. The model consists of $N$ RTPs, of diameter $\sigma$, on a lattice of $n=V/\sigma\in\mathbb{Z}^{+}$ sites. In this model the run durations $k_{i}$, for particle $i$, are exponentially distributed with rate-parameter $K^{-1}\geq 0$ and, during the runs, the particles have a constant self-propulsion speed $v_{0}$[21]. At passivity ($K\rightarrow 0$) the model reduces to the simple symmetric exclusion process (SSEP) [22]. The RTP particle model can be mapped to a mass transfer model with fluctuating directed bonds (as is done in the Dandekar-Chakraborti-Rajesh (DCR) model [23]), which we exploit here. The neighboring probability $P_{C}$ can be broken down as $2P_{C}={p_{\rightarrow,\leftarrow}+2p_{\rightarrow,\rightarrow}+p_{\leftarrow,\rightarrow}},$ where $p_{\rightarrow,\leftarrow}$ is the probability that two particles are next to each other with the left one moving right and the right one moving left, $p_{\rightarrow,\rightarrow}\equiv p_{\leftarrow,\leftarrow}$ is the probability that two particles are next to each other where both particles move in the same direction, and $p_{\leftarrow,\rightarrow}$ is the probability to be next to each other with the left particle moving left and the right particle moving right. The expressions for $_{\rightarrow,\leftarrow}, p,p_{\rightarrow,\rightarrow}$ and $p_{\leftarrow,\rightarrow}$ are difficult to obtain exactly, and so we resort to a mean-field approximation (14) (as is done in [23]). For weak activity, i.e., $K\lesssim 1$, the mean-field $P_{C}$ reads (up to $O(K^{3})$) as,
$P_{C}$ $\underset{K<<1}{=}\eta+\frac{\eta(\eta-1)^{2}K}{2}+\frac{(\eta-1)^{2}\eta(4\eta^{2}-7\eta+1)K^{2}}{8}.$ (8)
@@
The non-dimensionalized averaged cavity size (3) in the presence of weak and strong activity is then, respectively,
@@
$\frac{\langle N_{\Delta}\rangle}{N}$ $\underset{K\lesssim 1}{=}1-\eta-\frac{K(\eta-1)^{2}\eta}{2}$ (10)
$\qquad-\frac{K^{2}(\eta-1)^{2}\eta(1-7\eta+4\eta^{2})}{8},$
@@
$\frac{\langle V_{I}\rangle}{V}$ $\underset{K\lesssim 1}{=}(1-\eta)^{2}+\frac{K(\eta-1)^{2}\eta}{2},$ (11)
$\qquad+\frac{K^{2}(\eta-1)^{2}\eta(1-7\eta+4\eta^{2})}{8},$
@@
To test our theoretical predictions (9-11) we perform on- and off-lattice numerical simulations of repulsive RTPs in $d=1$ (Fig. 2). The implementation of the on-lattice model consists of RTPs hopping between adjacent sites and tumbling at exponentially-sampled rates with rate parameter $K^{-1}$, with a successful move to a new site only if that site is unoccupied. The off-lattice model has essentially the same ingredients, with the exception of continuous space and strict volume-exclusion being modeled by a repulsive (WCA) potential (16) (see Appendix B). In the simulations, for a given configuration $\omega$ the IS can be quantified through the insertion profile (Fig. D.1) which reads as:
@@
$\frac{N_{\Delta}(\omega)}{N}=\frac{1}{N}\sum_{i=1}^{N}\Theta(\tilde{\Delta}_{i,i+1}(\omega)),$ $\quad\frac{V_{I}(\omega)}{V}=\frac{1}{V}\int\limits_{0}^{V}\rho_{I}(x)dx,$
@@
Thus far, we have only considered one spatial dimension. We now determine how the insertion space changes for repulsive active particles in $d=2$. Analytical results in higher dimensions present significant challenges, even in equilibrium. The mathematical challenge is that, unlike in $d=1$, different particles can form the boundary of a single cavity and thus cavities cannot be labeled by particle indices consistently between different configurations. One certain difference in higher dimensions is that the number of cavities per particle approaches (in the dilute limit) $\langle N_{\Delta}\rangle\rightarrow 1$, in contrast with $\langle N_{\Delta}\rangle\rightarrow N$ for $d=1$ [16]. Another fact is that $\langle S_{I}\rangle$, the surface area of the insertion volume, is not a priori a constant.
@@ Appendix B Off-lattice exponential-tumble rate model of run-and-tumble particles (RTPs) in $d$ dimensions
where $\tau_{i}$ is an internal clock variable which is initially a uniformly distributed variable, $\mathbf{\hat{u}}_{i}(t)$ is a vector uniformly distributed on the $d$-dimensional sphere at time $t$, $s_{i}$ is the last time particle $i$ tumbled. The last two terms of the dynamics (16) impose the running, when $\tau_{i}$ is not a whole multiple of the $k_{i}$, and tumbling, otherwise. The passive (Markovian) limit is reached when $K\rightarrow\delta t$, where $\delta t$ is the lowest time resolution (0 in the continuous time limit), ensuring uniformly distributed tumbles at every time increment.
@@ Appendix C Off-lattice model of repulsive active Ornstein-Uhlenbeck particles (AOUPs) in $d$ dimensions
where $\dot{\mathbf{r}}_{i}$ is the instantaneous velocity of particle $i$, $\mu$ is the mobility, $D$ is the translational diffusion coefficient, $\mathbf{v}_{i}$ is the self propulsion governed by an Ornstein-Uhlenbeck process, $\tau$ is the persistence time, $D^{\prime}$ is the amplitude of the active noise, and $\mathbf{W}_{i}$ is an independent gaussian white noise process.
@@ Appendix D Numerical simulations for the off-lattice models
In the numerical simulations of the off-lattice models, the dynamics (16) and (17) are numerically solved using the Euler-Marayama method with timestep $\delta t=0.002$. Initially the particles are distributed uniformly in a volume larger than the chosen $V$ ($V^{\prime}>V(\eta)$) and then a pre-simulation is performed to drag the particles towards the origin until each particle is within $V(\eta)$ and they are then made to relax according to equilibrium dynamics. For all simulations we impose periodic boundary conditions. Production runs begin, at the chosen activity parameter, once the distribution of particles is sufficiently uniform. We gather IS statistics every 1000 timesteps and our minimum threshold sample size is $N_{s}=5000$ at each $K$, or $\tau$, and $\eta$.
@@ Appendix E $\langle N_{\Delta}\Delta\rangle=\langle N_{\Delta}\rangle\langle\Delta\rangle$?
Figure E.1: Plot showing the growing statistical dependence of $N_{\Delta}$ and $\bar{\Delta}$ for increased activity (higher $K$) for off-lattice (top) and off-lattice (bottom) simulations of RTPs in $d=1$. Close to equilibrium $K=1$ statistical independence is a valid assumption (besides very high packing fraction in the lattice data).
@@
$\langle V_{I}\rangle=\langle N_{\Delta}\bar{\Delta}\rangle=\langle N_{\Delta}\rangle\langle\bar{\Delta}\rangle_{\omega}+\mathopen{\hbox{\set@color${\langle}$}\kern-1.94444pt\leavevmode\hbox{\set@color${\langle}$}}N_{\Delta}\times\bar{\Delta}\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\leavevmode\hbox{\set@color${\rangle}$}},$ (20)
where $\mathopen{\hbox{\set@color${\langle}$}\kern-1.94444pt\leavevmode\hbox{\set@color${\langle}$}}N_{\Delta}\times\bar{\Delta}\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\leavevmode\hbox{\set@color${\rangle}$}}:=\langle N_{\Delta}\bar{\Delta}\rangle-\langle N_{\Delta}\rangle\langle\bar{\Delta}\rangle_{\omega}$ is the two-point connected correlation function. In equilibrium one can get away with $\langle N_{\Delta}\bar{\Delta}\rangle=\langle N_{\Delta}\rangle\langle\bar{\Delta}\rangle_{\omega}$ due to the statistical independence of the number of cavities and the averaged (per configuration) cavity size (see Fig. E.1).
@@ Appendix F Method to determine the IS in $d=2$
      Form the particle-centered bounding interval in cell units: compute integer bounds $n_{x-\sigma},n_{x+\sigma},n_{y-\sigma},n_{y+\sigma}$ (see Fig. F.1).

### v2

@@ Abstract
where $\beta=(k_{B}T)^{-1}$ is the inverse temperature, $P$ is the bulk pressure, $V$ is the total volume of the space, $N$ is the number of constituents, $\sigma$ is the particle diameter (or repulsion range), $d$ is the spatial dimension, $\Lambda$ is the thermal de Broglie wavelength, $\langle V_{I}\rangle$ is the averaged (over all values in all configurations) total insertion volume, and $\langle S_{I}\rangle$ its averaged concomitant surface. The insertion volume is expressed in terms of the averaged (over all values across all configurations) number of cavities and cavity size as $\langle V_{I}\rangle=\langle N_{\Delta}\rangle\langle\Delta\rangle$ [15, 17].
@@
Figure 2: Insertion space (IS) for repulsive active particles in $d=1$. a) Averaged normalized cavity size as a function of packing fraction $\eta$, as determined from off-lattice (filled points) and on-lattice (unfilled points) simulations of RTPs at various $K$. Analytical results are solid lines; for $K\leq 10.0$ the lines are the weak activity mean-field expressions while for $K>10.0$ we show the high-activity expressions (9-11). Exact passive expressions are from (7). b) The averaged total insertion volume divided by the system volume. c) Averaged number of insertion cavities per particle as a function of $\eta$. d) The geometric compressibility, $Z_{G}$, as a function of $\eta$. Inset is the geometric chemical potential $\mu_{G}$ against $\eta$. e) - h) The same quantities as for a)-d) but for the AOUP (off-lattice) model. High-activity analytics (for $\tau=1000$) are (9-11) with $K\simeq\sqrt{D^{\prime}\tau}/\sigma$. Parameters (RTPs): $N=1024$, $\mu=1$ (off-lattice), $\epsilon=10$ (off-lattice), and $\sigma=1$; AOUPs: $N=1024$, $\mu=1$, $\epsilon=10$, $D^{\prime}=\frac{1}{2}$ and $\sigma=1$.
Firstly, in $d=1$, the dimensionless averaged cavity size $\left\langle\Delta\right\rangle/\sigma\equiv\langle\tilde{\Delta}\rangle$ is the free (unoccupied) space divided by the number of non-touching particle pairs minus one and reads as:
@@
where $P_{C}$ is the probability of two particles being in contact and $\eta=N\sigma/V$ is the packing fraction. In one dimension, $\langle S_{I}\rangle=2\langle N_{\Delta}\rangle$ exactly, implying that the geometric compressibility $Z_{G}$ (1) can be simplified as [13]
@@
with $\langle V_{I}\rangle=V-N(2\sigma)$ in the dilute case, where $2\sigma$ is the volume excluded by one particle, and $\langle V_{I}\rangle=0$ at full packing. Note, that $P_{C}$ is absent in the original expressions of these quantities [13]. The equilibrium one-dimensional hard-sphere model is the Tonks gas, which is exactly solvable [8], where the IS quantities are given exactly as [14, 15]:
@@
For the active model we consider repulsive run-and-tumble particles (RTPs) in $d=1$. The model consists of $N$ RTPs, of diameter $\sigma$, on a lattice of $n=V/\sigma\in\mathbb{Z}^{+}$ sites. In this model the run durations $k_{i}$, for particle $i$, are exponentially distributed with rate-parameter $K^{-1}\geq 0$ and, during the runs, the particles have a constant self-propulsion speed $v_{0}$[21]. At passivity ($K\rightarrow 0$) the model reduces to the simple symmetric exclusion process (SSEP) [22]. The RTP particle model can be mapped to a mass transfer model with fluctuating directed bonds (as is done in the Dandekar-Chakraborti-Rajesh (DCR) model [23]), which we exploit here. The neighboring probability $P_{C}$ can be broken down as $4P_{C}={p_{\rightarrow,\leftarrow}+2p_{\rightarrow,\rightarrow}+p_{\leftarrow,\rightarrow}},$ where $p_{\rightarrow,\leftarrow}$ is the probability that two particles are next to each other with the left one moving right and the right one moving left, $p_{\rightarrow,\rightarrow}\equiv p_{\leftarrow,\leftarrow}$ is the probability that two particles are next to each other where both particles move in the same direction, and $p_{\leftarrow,\rightarrow}$ is the probability to be next to each other with the left particle moving left and the right particle moving right. The expressions for $_{\rightarrow,\leftarrow}, p,p_{\rightarrow,\rightarrow}$ and $p_{\leftarrow,\rightarrow}$ are difficult to obtain exactly, and so we resort to a mean-field approximation (14) (as is done in [23]). For weak activity, i.e., $K\lesssim 1$, the mean-field $P_{C}$ reads (up to $O(K^{3})$) as,
$P_{C}$ $\underset{K\lesssim 1}{=}\eta+\frac{\eta(\eta-1)^{2}K}{2}+\frac{(\eta-1)^{2}\eta(4\eta^{2}-7\eta+1)K^{2}}{8}.$ (8)
@@
The dimensionless averaged cavity size (3) in the presence of weak and strong activity is then, respectively,
@@
$\frac{\langle N_{\Delta}\rangle}{N}$ $\underset{K\lesssim 1}{=}(1-\eta)-\frac{K\eta(1-\eta)^{2}}{2}\left(1-\frac{K(1-7\eta+4\eta^{2})}{4}\right),$ (10)
@@
$\frac{\langle V_{I}\rangle}{V}$ $\underset{K\lesssim 1}{=}(1-\eta)^{2}\bigg(1+\frac{K\eta}{2}+\frac{K^{2}\eta(1-7\eta+4\eta^{2})}{8}\bigg),$ (11)
@@
To test our theoretical predictions (9-11) we perform on- and off-lattice numerical simulations of repulsive RTPs in $d=1$ (Fig. 2). The implementation of the on-lattice model consists of RTPs hopping between adjacent sites and tumbling at exponentially-sampled rates with rate parameter $K^{-1}$, with a successful move to a new site only if that site is unoccupied. The off-lattice model has essentially the same ingredients, with the exception of continuous space and strict volume-exclusion being modeled by a repulsive (WCA) potential (see the dynamics (16) and Appendix B). In the simulations, for a given configuration $\omega$ the IS can be quantified through the insertion profile (Fig. D.1) which reads as:
@@
$\frac{N_{\Delta}(\omega)}{N}=\frac{1}{N}\sum_{i=1}^{N}\Theta(\tilde{\Delta}_{i,i+1}(\omega)),$ $\quad\frac{V_{I}(\omega)}{V_{d=1}}=\frac{1}{V}\int\limits_{0}^{V}\rho_{I}(x)dx,$
@@
Thus far, we have only considered one spatial dimension. We now determine how the insertion space changes for repulsive active particles in $d=2$. Analytical results in higher dimensions present significant challenges, even in equilibrium. The mathematical challenge is that, unlike in $d=1$, different particles can form the boundary of a single cavity and thus cavities cannot be labeled by particle indices consistently between different configurations. One certain difference in higher dimensions is that the number of cavities per particle approaches (in the dilute limit) $\langle N_{\Delta}\rangle\rightarrow 1$, in contrast with $\langle N_{\Delta}\rangle\rightarrow N$ for $d=1$ [16]. Another fact is that $\langle S_{I}\rangle/\langle N_{\Delta}\rangle$, the typical surface area of the insertion cavity, is not a priori a constant.
@@ Appendix B Off-lattice exponential-tumble rate model of run-and-tumble particles (RTPs) in $d$ dimensions
where $\tau_{i}$ is an internal clock variable which is initially a uniformly distributed variable, $\phi(\{\mathbf{r}\})$ is the total potential energy imposing the repulsion between particles and is defined as:
$\phi=\frac{1}{2}\sum_{i,j}^{N}U(r_{ij}),$ (17)
where $r_{ij}\equiv|\mathbf{r}_{i}-\mathbf{r}_{j}|$ and $U$ is the interaction pair potential given in WCA form:
$U(r)=\left[4\varepsilon\left(\left(\frac{\sigma}{r}\right)^{12}-\left(\frac{\sigma}{r}\right)^{6}\right)+\varepsilon\right]\Theta(2^{1/6}\sigma-r),$ (18)
where $\varepsilon$ is the repulsion strength, $\sigma$ is the interaction range. The quantity $\mathbf{\hat{u}}_{i}(t)$ in the dynamics is a vector uniformly distributed on the $d$-dimensional sphere at time $t$, $s_{i}$ is the last time particle $i$ tumbled. The last two terms of the dynamics (16) impose the running, when $\tau_{i}$ is not a whole multiple of the $k_{i}$, and tumbling, otherwise. The passive (Markovian) limit is reached when $K\rightarrow 0$, ensuring uniformly distributed tumbles at every time increment.
@@ Appendix C Off-lattice model of repulsive active Ornstein-Uhlenbeck particles (AOUPs) in $d$ dimensions
where $\dot{\mathbf{r}}_{i}$ is the instantaneous velocity of particle $i$, $\mu$ is the mobility, $\mathbf{v}_{i}$ is the self propulsion governed by an Ornstein-Uhlenbeck process, $\tau$ is the persistence time, $D^{\prime}$ is the amplitude of the active noise, and $\mathbf{W}_{i}$ is an independent gaussian white noise process.
@@ Appendix D Numerical simulations for the off-lattice models
In the numerical simulations of the off-lattice models, the dynamics (16) and (19) are numerically solved using the Euler-Maruyama method with timestep $\delta t=0.002$. Initially the particles are distributed uniformly in a volume larger than the chosen $V$ ($V^{\prime}>V(\eta)$) and then a pre-simulation is performed to drag the particles towards the origin until each particle is within $V(\eta)$ and they are then made to relax according to equilibrium dynamics. For all simulations we impose periodic boundary conditions. Production runs begin, at the chosen activity parameter, once the distribution of particles is sufficiently uniform. We gather IS statistics every 1000 timesteps and our minimum threshold sample size is $N_{s}=5000$ at each $K$, or $\tau$, and $\eta$.
@@ Appendix E $\langle N_{\Delta}\Delta\rangle=\langle N_{\Delta}\rangle\langle\Delta\rangle$?
Figure E.1: Plot showing the growing statistical dependence of $N_{\Delta}$ and $\bar{\Delta}$ for increased activity (higher $K$) for off-lattice (top) and on-lattice (bottom) simulations of RTPs in $d=1$. Close to equilibrium $K=1$ statistical independence is a valid assumption (besides very high packing fraction in the lattice data).
@@
$\langle V_{I}\rangle=\langle N_{\Delta}\bar{\Delta}\rangle=\langle N_{\Delta}\rangle\langle\bar{\Delta}\rangle_{\omega}+\mathopen{\hbox{\set@color${\langle}$}\kern-1.94444pt\hbox{\set@color${\langle}$}}N_{\Delta}\times\bar{\Delta}\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\hbox{\set@color${\rangle}$}},$ (22)
where $\mathopen{\hbox{\set@color${\langle}$}\kern-1.94444pt\hbox{\set@color${\langle}$}}N_{\Delta}\times\bar{\Delta}\mathclose{\hbox{\set@color${\rangle}$}\kern-1.94444pt\hbox{\set@color${\rangle}$}}:=\langle N_{\Delta}\bar{\Delta}\rangle-\langle N_{\Delta}\rangle\langle\bar{\Delta}\rangle_{\omega}$ is the two-point connected correlation function. In equilibrium one can get away with $\langle N_{\Delta}\bar{\Delta}\rangle=\langle N_{\Delta}\rangle\langle\bar{\Delta}\rangle_{\omega}$ due to the statistical independence of the number of cavities and the averaged (per configuration) cavity size (see Fig. E.1).
@@ Appendix F Method to determine the IS in $d=2$
      Form the particle-centered bounding interval in cell units: compute integer bounds $n_{x-\sigma},n_{x+\sigma},n_{y-\sigma},n_{y+\sigma}$ (see Fig. F.1). An example of an integer bounds can be computed as $n_{x-\sigma}=\left\lceil((x_{0}-\sigma)-(-L_{x}/2))/l\right\rceil$.